OpenStudy (anonymous):

Which lines passes through the point (0,3) and has a slope of -1/2? A. y=-2x-3 B. y=-1/2x+3 C. y=1/2x-3 D. y=-1/2x-3/2 E. y=-2x-3/2

OpenStudy (anonymous):

@jojo4eva

OpenStudy (anonymous):

y = mx + b when m = slope and b = y-intercept. -1/2 = m and 3 = b So your equation is y = -1/2x + 3